### v2

# arXiv:2602.21068v2[stat.ME] 27 Feb 2026

Abstract
Experimental evaluations of public policies often randomize a new intervention within many sites or blocks. After a report of an overall result — statistically significant or not — the natural question from a policy maker is: where did any effects occur? Standard adjustments for multiple testing provide little power to answer this question. In simulations modeled after a 44-block education trial, the Hommel adjustment — among the most powerful procedures controlling the family-wise error rate (FWER) — detects effects in only 11% of truly non-null blocks. We develop a procedure that tests hypotheses top-down through a tree: test the overall null at the root, then groups of blocks, then individual blocks, stopping any branch where the null is not rejected. In the same 44-block design, this approach detects effects in 44% of non-null blocks — roughly four times the detection rate. A stopping rule and valid tests at each node suffice for weak FWER control. We show that the strong-sense FWER depends on how rejection probabilities accumulate along paths through the tree. This yields a diagnostic: when power decays fast enough relative to branching, no adjustment is needed; otherwise, an adaptive $\alpha$-adjustment restores control. We apply the method to 25 MDRC education trials and provide an R package, manytestsr.
1 Where do block-level effects occur?
Consider a block-randomized experiment with 44 experimental blocks such as the Detroit Promise Program [undefr] where treatment was assigned to students of different academic cohorts by five community colleges. The overall test rejects the null of no effects. The policy-maker considering changes across all of Detroit’s community colleges now asks: where did the effects occur? Were they concentrated in a few sites, or spread evenly? The answer determines whether to replicate the intervention everywhere or investigate why it succeeded in some places and failed in others.
A natural statistical response — test the null hypothesis of no effects in each block separately, then adjust for multiplicity — is disappointing. In data we simulated to follow the design of the Detroit Promise Program, the Hommel adjustment, one of the most powerful procedures that controls the family-wise error rate (FWER), detects effects in only 11% of the blocks that truly have them. The analyst reports “not enough information,” and the policy-maker gets nothing useful.
This paper presents a different approach. [^1] We focus on the family-wise error rate (FWER) rather than the false discovery rate (FDR). The FDR controls the expected proportion of false discoveries among rejections, which suits settings where rejected hypotheses will be validated in follow-up studies. The FWER controls the probability of any false rejection, which suits settings where rejected hypotheses lead directly to policy action. When a school district decides to scale an intervention based on where effects were detected, a single erroneous claim can misdirect resources. FDR-oriented extensions of tree-structured testing are discussed in Section 5. Instead of testing every block and adjusting afterward — a bottom-up strategy — we test hypotheses top-down on a tree. We begin with the overall null at the root. If rejected, we split the blocks into groups following the administrative structure of the experiment and test each group. If a group’s null is rejected, we split again and continue down to individual blocks. If a group’s null is not rejected, we stop testing on that branch. In the same simulation, this structured approach detects effects in 44% of the truly non-null blocks while controlling the FWER at the nominal level — roughly four times the detection rate of the bottom-up alternative.
Two conditions drive this result. First, a stopping rule: test a child hypothesis only after rejecting its parent. Second, valid tests: each individual test controls its false positive rate at level $\alpha$. We show that two conditions suffice for weak FWER control — control when all null hypotheses are true. For strong control, we analyze how power decay from data splitting limits error accumulation through the tree and, when needed, apply an adaptive $\alpha$-adjustment. This framework builds on the closed testing tradition of [undefo], [undeft] and the sequential structured testing of [undefe], [undefd], and [undefp], which we extend to randomization-based inference in block-randomized experiments.
Our extensions to this foundational literature are substantive rather than cosmetic. In genomic applications of tree-structured testing, the same data are analyzed at every node, the analyst must impose the conditions for valid testing as assumptions, and FWER adjustments apply uniformly across the tree. Block-randomized experiments differ in ways that the existing framework does not exploit. First, the experimental design guarantees valid tests at each node: randomization-based $p$-values have exact size control by construction. Second, testing subsets of blocks is itself a form of data splitting — child tests use less data and so should not produce more extreme evidence than their parents — so power decays through the tree as sample sizes shrink at each split, creating a geometric damping of error along each path. We derive an explicit expression for FWER as a function of these path-wise rejection probabilities (Proposition 1), which reveals that high power at the root can inflate error rates at the next level while low power at deep levels provides natural protection from false positive errors. This expression motivates an adaptive $\alpha$-adjustment that is stringent where power is high and relaxed where power decay triggers the stopping rule and testing stops. This paper makes three contributions. First, we show that the conditions for tree-structured weak FWER control are satisfied by design in block-randomized experiments. Second, we derive the relationship between power decay and strong FWER that makes adaptive adjustment possible. Third, we demonstrate that these ideas yield two- to six-fold more discoveries than standard adjustments in realistic multi-site trials (Table 5), with gains of an order of magnitude or more in wider trees (Table 4). For researchers working in policy evaluation, the practical implication is clear: in experiments with administrative structures, more can be learned about the welfare effects of a new policy than we might have imagined without testing hypotheses in order.
Two traditions of research address heterogeneous treatment effects. One estimates conditional average treatment effects as functions of observed covariates [undefab, undeff, e.g.]. The other studies the distribution of individual effects directly, asking how many units benefited and by how much [undefh, undefm, e.g.]. Our work is closer to the second tradition: rather than asking which covariates moderate effects, we ask which experimental blocks or groups of blocks show detectable effects. Where covariate-based approaches focus on columns of the data matrix, we focus on rows — the sites, cohorts, and schools where effects did or did not occur. The approaches are complementary; detections could guide subsequent covariate-based investigation of why effects vary. A related question arises when sites differ not only in treatment effects but in local ecological conditions that confound between-site comparisons. [undefk] develop methods to isolate organizational effectiveness from such confounding in multisite trials using parametric models and empirical bayes style statistical inference.
We focus on block-randomized experiments with two treatment arms and roughly continuous outcomes using randomization-based statistical inference. The results generalize to binary outcomes and multi-arm trials directly, although we do not engage those designs in our simulations or application.
The paper proceeds as follows. Section 2 develops the bottom-up and top-down testing approaches, proves weak FWER control (Theorem 2), and illustrates the performance of the procedure using simulations across a wide range of tree sizes. Section 3 addresses strong FWER control by analyzing how power decay through the tree naturally limits error accumulation, and develops an adaptive $\alpha$-adjustment for settings where this natural gating is insufficient. We then demonstrate the method on data simulated to follow the Detroit Promise Program design (Section 3.2) and apply it to 25 block-randomized education trials fielded by the MDRC [undefa]. [^2] In the supplement we develop a test statistic sensitive to distributional differences beyond mean shifts, so that effects that are positive in some blocks and negative in others do not cancel at the root. This statistic draws on energy statistics [undefz] and multivariate permutation tests [undefx, undefl].
2 Testing to Detect Effects in Blocks
Block-randomized experiments allow two styles of testing procedures that target block-specific causal effects. In a study like that shown in Figure 1, we have a set of blocks, $B$, within which a new policy intervention is randomly assigned to $m_{b}$ people, leaving $n_{b}-m_{b}$ in the status quo. Notice the inverted tree-like shape of the experiment. The individual blocks are on the bottom of the tree and, in this case, are nested within administrative units like cities and states toward the top of the tree.
{forest}
Figure 1: An administratively organized structure of blocks. A study randomly assigns people within offices ($B$) to a new intervention. Each office is an experimental block containing $m_{b}$ people assigned to the intervention and $n_{b}-m_{b}$ people assigned to the status quo.
2.1 Testing in every block: the basic problem of multiple testing
Since each block contains multiple units assigned at random to treatment, an analyst could treat each block as a mini-experiment [undefc]. With 100 blocks, we could execute 100 tests of the null hypothesis of no treatment effects. We should also adjust the rejection criteria for those tests to protect the decision-makers from making too many false positive errors. Why adjust? Imagine that, in fact, there were no effects in any of the 100 blocks. If the tests are independent of one another, and we have set our rejection threshold at $\alpha=0.05$, then we know that the probability that at least one of the 100 unadjusted tests would yield a statistically significant result, $p<0.05$, in error is: $\mathbb{P}(\text{at least one }p<0.05)=(1-(1-0.05)^{100})\approx 0.99$. That is, if we did a test in each of the 100 blocks in an experiment in which no effects occurred in any block, we would virtually certainly reject this null in one or more blocks — and make an error when we interpret this rejection of "no effects" in those blocks as implying some true causal effect occurred in those blocks. This error rate of a collection of tests is often called the Family-wise Error Rate (FWER), and it can be controlled by an adjustment. For example, the Bonferroni adjustment suggests that if we refuse to reject the null hypothesis unless $\alpha_{\text{adjusted}}=p\leq 0.05/100=0.0005$, then we will have controlled the FWER because P(at least one p<0.05)=(1(10.05/100)100)0.05P($\text${at least one }p<0.05)=$(1-$$^{100})\approx 0.05$ [undefn]. [^3] Technically speaking, this explanation shows that dividing the rejection threshold by the number of tests, controls the FWER in the “weak sense”, which means it controls the FWER when there are no effects in any block. We will discuss strong and weak control of the FWER later in the paper.
Of course, this approach makes it much harder to detect effects when they are actually present. For example, the power of a t-test to detect an effect of 0.8 standard deviations within a block of 50 units is 0.79 when $\alpha=0.05$ and nearly half that power (power=0.46) when $\alpha=0.005$. Structured approaches to testing can increase this power while controlling the FWER.
2.2 Sequential Structured Testing and Weak Control of the FWER
Figure 1 shows that each experimental block is nested within a hierarchy. When hypotheses are nested in this way, they can be tested in a specific order under rules that avoid the severe power loss of bottom-up adjustments. This insight originates in the “closed testing” framework of [undefo], developed for structured experiments by [undeft, undefw]. We build especially on the general rules for sequential nested testing articulated by [undefe, undefd] and [undefp]. We explain the reasoning behind the structured testing approach here before turning to formal development, simulations, and application.
The general idea is to start testing hypotheses at the “root” of the tree — testing first the null hypothesis of no effects in any block. If we can reject that hypothesis then we know either (1) that at least one block contains at least one unit with a non-zero treatment effect or (2) that we have just made a false positive error. This first test could be the same test that anyone analyzing a block-randomized experiment might use — for example a block-aligned t-test or rank-based Wilcoxon test. [^4] We discuss test statistics in Supplement C. This project led us to develop two approaches to assessing the hypothesis of no effects that are immune from positive versus negative effect cancelation problems. We defer the test-statistic development to the Supplement and focus here on the tree-structured testing procedure. The second and subsequent tests occur “lower” in the tree on subsets of the data split into disjoint groups. Once a test does not reject a hypothesis, testing stops on that branch of the tree. We prove that this procedure controls the FWER when all null hypotheses are true (Theorem 2) and confirm the result via simulation across a wide range of tree sizes (Table 1).
Conditions for FWER control
To fix ideas, consider the tree in Figure 2. Write $H_{i}$ for the null hypothesis of no treatment effects among any of the units in the blocks descended from node $i$. At the root, $H_{1}$ is the hypothesis that no unit in any block has a treatment effect. At a leaf — nodes 5 through 13 in the figure — $H_{i}$ pertains to a single experimental block. We write $p_{i}$ for the $p$-value from the test of $H_{i}$.
$\text{node }1$$\text{node }2$$\text{node }3$$\text{node }4$}5{5}6{6}7{7}8{8}9{9}10{10}11{11}12{12}13{13}
Figure 2: A $k$-ary tree with $k=3$ nodes per level and $L=3$ levels and $k^{L-1}=9$ terminal nodes or “leaves" representing individual experimental blocks.
We state two conditions. Together they suffice for weak FWER control (Theorem 2). Combined with an analysis of power decay and, when needed, adaptive $\alpha$-allocation, they also provide strong control (§ 3).
Condition 1 (Stopping rule).
A hypothesis $H_{i}$ is tested only if every ancestor of node $i$ has been rejected. If the root hypothesis is not rejected, no further tests are performed. If any hypothesis on a path from root to leaf is not rejected, no descendant on that branch is tested at all. In Figure 2, we do not test $H_{5}$ unless we have rejected both $H_{2}$ and $H_{1}$.
This condition distinguishes our procedure from tree-structured testing frameworks that test every node and then determine which rejections to keep [undefe, undefd] and connects with frameworks for strictly nested testing that would focus on a single branch of a tree [undeft]. Our procedure may test only a small fraction of the tree’s nodes — and each untested node is a test that cannot produce a false positive.
Condition 2 (Valid tests).
Each test at a node, if executed on its own, has a false positive rate no greater than $\alpha$. In block-randomized experiments, randomization-based tests satisfy this by design: permuting the treatment assignment within blocks produces $p$-values with guaranteed size control [undefu, Chapter 2].
These conditions enable us to state and prove the following theorem:
{theorem}
[] Conditions 1 and 2 suffice for weak FWER control
A family of true null hypotheses organized on an irregular or regular $k$-ary tree and tested following the stopping rule (Condition 1) with valid tests at each node (Condition 2) will produce a family-wise error rate (FWER) no greater than $\alpha$. We call a $k$-ary tree “regular” if the number of child nodes of a parent is the same for all nodes in the tree. A $k$-ary tree is “irregular” if the number of child nodes of a parent node may differ within and across levels of the tree.
The proof of Theorem 2 is in Appendix A in the Supplement. We provide some intuition here.
Figure 3 sketches the algorithm that follows those rules while adding the idea of splitting the blocks into groups. Data splitting will be important when we engage with strong control of the FWER, but is not necessary for weak control as we show in the proof and in the simulations below. In this version we talk about “splitting” the blocks as a way to speak in general about testing hypotheses on subsets of blocks: we do not propose to split within blocks, just to divide the experimental blocks into groups of blocks. In the example of the administrative structure above the “splits” are fixed by design.
In Figure 3 we write $\mathcal{B}_{1}$ to refer to the set of all the experimental blocks in the study — the set of blocks containing the units relevant to the hypothesis at node 1, $H_{1}$ — and $\mathcal{B}_{2}$ and $\mathcal{B}_{3}$ for mutually exclusive subsets of blocks nested within $\mathcal{B}_{1}$ (which could be individual blocks that cannot be further split or groups of blocks which could be divided). We write $\mathcal{B}_{4}$ for the first subset of the $\mathcal{B}_{2}$ set and so on. For simplicity in this paper we only test the null hypothesis of no effects. At the level 1 or the root level, we test the null of no effects for any unit in any block and produce a $p$-value as the output of this test that we call $p_{\text{node }1}\equiv p_{1}$ for “$p$-value for the node 1 hypothesis” or “$p$-value for the hypothesis that the potential outcomes for all units in all blocks under treatment are the same as the potential outcomes under control, $H_{0}:y_{i,b,1}=y_{i,b,0}\;\forall i,b\in\mathcal{B}_{1}$.” If we cannot reject this overall null, such that $p_{1}>\alpha$, then testing stops. If we can reject the overall null, then at least one unit in one block has a causal effect, we split the data into subsets and do further tests as shown in Figure 3.
[…]
Figure 3: Simplified flow of the Top-Down Testing and Splitting Algorithm with fixed false positive level $\alpha$. All blocks are in set $\mathcal{B}_{1}$, $\mathcal{B}_{2}$ is a subset of $\mathcal{B}_{1}$, $\mathcal{B}_{4}$ is a subset of the blocks in $\mathcal{B}_{2}$. The $p$-value, $p_{1}$, is the result from a test of the hypothesis of no effects using all the blocks (i.e using the set $\mathcal{B}_{1}$), $p_{2}$ is the $p$-value from a test of the null of no effects using only the blocks in $\mathcal{B}_{2}$. Testing stops when $p>\alpha$ or when the number of blocks in $\mathcal{B}$, written $|\mathcal{B}|$, is 1 such that for a given node $i$, $|\mathcal{B}_{i}|=1$.
The intuition for the proof of Theorem 2 arises from the fact that in a randomized experiment, we know that the probability of a false positive error in a single randomization-based test is low and controlled: the size is less than or equal to the level ($\alpha$). This is stated as a condition of the proof in Condition 2. Another way to say this: if the null of no effects is true in all blocks, then we should reject the root hypothesis with probability less than $\alpha$. An incorrect rejection thus happens with very low probability and so rarely are other tests even done, and when they are done their own false positive rates are less than or equal to $\alpha$. The simulation results in Table 1 show control of the FWER for a wide range of tree sizes and configurations when there are no treatment effects.
2.2.1 Simulation Study of Weak Control of the FWER to Illustrate the Proof
Given a specification of a complete $k$-ary tree using $k$ nodes per level and $L$ levels we drew $p$-values from uniform distributions. The valid tests condition (Condition 2) implies that the distribution of $p$-values across tests of the true null should be stochastically dominated by the uniform [undefn, Chapter 3]. That is, in many tests of a true null hypothesis, only 5% of $p$-values should be less than $p=0.05$, 10% should be less than $p=0.1$, etc….
  1. 1.
    For the root node, draw $p_{1}\sim U(0,1)$; a draw which respects the valid test condition such that $\mathbb{P}(p_{1}\leq\alpha)\leq\alpha$ in this case.
  2. 2.
    If $p_{1}\leq\alpha$, draw $p_{l=2,k}$ independent $p$-values from $U(0,1)$ for the $k$-nodes at level 2.
  3. 3.
    For any of the $k$ $p$-values at level 2 where $p_{l=2,k}\leq\alpha$, generate $p$-values for its children with $U(0,1)$. For any $p_{l=2,k}>\alpha$, stop testing in the branch of the tree that would descend from that node. This implements the stopping rule condition (Condition 1).
  4. 4.
    Continue testing, drawing from uniform distributions for child nodes with parents with $p\leq\alpha$, descend into the tree towards the leaves, and stop testing in any branch when $p>\alpha$ or when the node tested is a leaf.
In this simulation we did not represent the idea of data splitting or any other adjustments. So, these results are much more optimistic than they would be if, in fact, the sample size used to test a hypothesis at level 2 is less than that available at the root node.
For each tree, we repeated this procedure 10,000 times, recording whether any of the $p$-values in the tree of true null hypotheses were $\leq\alpha$. The proportion of simulations with at least one such false rejection is a measure of the FWER.
Table 1 shows the results. Whether the tree has $k=2$ nodes per level or $k=100$, the rules lead to a maximum FWER within simulation error. [^5] Whether we use $2\times\sqrt{0.05(1-0.05)/10000}=0.004$ or $2\times\sqrt{0.5(1-0.5)/10000}=0.01$ we interpret these results to show nominal FWER control. A tree with $k$ nodes per level and a maximum of $L$ levels has $k^{\ell-1}$ nodes at level $\ell$ and $k^{L}$ leaves. The first row of the table shows binary trees ($k=2$): with 2 levels we have 4 leaves and 7 total nodes; with 18 levels we have 262,144 leaves and 524,287 total nodes. This set of simulations produces research designs that tend to be unrealistic from the perspective of block-randomized experiments in public policy — rarely would one have 260,000+ experimental blocks! But these extreme examples provide what we hope is a compelling illustration of Theorem 2.
Levels Max Avg. Max Nodes Leaves
$k$ Min Max FWER Tests Min Max Min Max
2 2 18 0.052 1.008 7 524287 4 262144
4 2 8 0.055 1.015 21 87381 16 65536
6 2 6 0.054 1.026 43 55987 36 46656
8 2 6 0.050 1.032 73 299593 64 262144
10 2 4 0.049 1.044 111 11111 100 10000
20 2 4 0.050 1.202 421 168421 400 160000
50 2 2 0.048 1.402 2551 2551 2500 2500
100 2 2 0.050 2.498 10101 10101 10000 10000
Table 1: This table shows weak-control of the family-wise error rate across 10,000 simulations (estimated simulation error = .01) using $\alpha=.05$ on $k$-ary trees with $k$ nodes per level where the hypothesis of no effects is true for all nodes. Each row summarizes the results of simulations for the trees with a given $k$ and between the minimum and maximum number of levels for that number of nodes per level. Maximum average false positive rates across the 10,000 simulations are shown in the ‘Max FWER’ column. The maximum number of nodes tested are shown in ‘Max Tests’. And information about the range of trees is also shown: the total number of nodes in the trees and the total number of terminal nodes or leaves.
Table 1 also shows that the two rules lead the algorithm to nearly always stop after testing the single root node: the average number of tests is close to 1, even when the tree has hundreds of thousands of nodes. In contrast, the bottom-up procedure must execute a test in every block and then adjust the results — thousands of tests where the top-down procedure does one [undefj].
2.2.2 Why might weak control suffice?
The simulations and proof above assume that no units in any blocks have treatment effects — that all null hypotheses are true. This is what it means for a procedure to control the FWER “in the weak sense” [undefi]. The next section develops conditions for strong FWER control, which holds regardless of which hypotheses are true. Yet Conditions 1 and 2 on their own may be appropriate for some research settings, and the reasoning is worth making explicit.
A useful point of comparison is the Benjamini–Hochberg procedure [undef], the standard tool for controlling the false discovery rate (FDR) — the expected proportion of false rejections among all rejections. [undef] proved that their procedure controls the FDR at level $\alpha$ regardless of which hypotheses are true and which are false. Yet when all null hypotheses are true, every rejection is a false rejection, so FDR equals FWER; the procedure therefore controls the FWER in the weak sense but not in the strong sense. Researchers accept this because FDR control suits exploratory settings where findings will be validated in follow-up studies. The tree-structured approach does not control the FDR (developing such guarantees is beyond the scope of this paper), but it shares with FDR procedures the property that the FWER is controlled at $\alpha$ when all null hypotheses are true.
The choice among error-rate guarantees depends on the research question. Does this program work at all? prioritizes detecting any effect over pinpointing which specific outcomes are affected; weak FWER control suits this goal. Because the gate absorbs the entire multiplicity cost, every downstream test runs at nominal $\alpha$ — power that both strong FWER and FDR methods sacrifice for their broader guarantees. Which individual outcomes are affected? — where each claimed detection must stand on its own — calls for strong FWER control. FDR procedures occupy a middle ground, accepting that the list of discoveries may contain false entries but limiting their proportion. For the first question, the tree-structured procedure is a natural fit: the overall test serves as a gate, and Conditions 1 and 2 make the weak FWER guarantee explicit.
Researchers who are using the tree-structured approach to screen blocks for closer examination can work with Conditions 1 and 2 alone. But in many applied settings, researchers want to make claims about specific blocks with the same error-rate guarantees they would demand of any confirmatory analysis. The next section shows how: analyzing the power decay inherent in data splitting — and, when needed, applying adaptive $\alpha$-allocation — yields strong FWER control while preserving the power advantage of top-down testing.
3 Strong control of the FWER
Conditions 1 and 2 control false positive errors when all null hypotheses are true. Control of the FWER when at least one block contains a non-zero treatment effect — strong control — requires additional structure. The key quantity turns out to be the power of the test at each node: the probability that the procedure both reaches a null node and then rejects it depends on the sequence of rejection probabilities along the path from root to that node.
To fix ideas, consider the small tree in Figure 4: a $k$-ary tree with $k=3$ children per node, $L=3$ levels. A complete $k$-ary tree has $k^{L-1}$ leaves, so this tree has $3^{2}=9$ leaves or experimental blocks and 13 total nodes. We develop our theory using complete $k$-ary trees for simplicity, though the applications use irregular trees. In the figure, the causal effect is non-zero only in block 5 (boxed). Because block 5 has a non-zero effect, its ancestors — nodes 2 and the root — are also non-null. So 3 of the 13 nodes are non-null and 10 are null.
123{3}4{4}56{6}7{7}8{8}9{9}10{10}11{11}12{12}13{13}
Figure 4: A $k$-ary tree with $k=3$ and $L=3$. Boxes show non-null nodes: since the leaf (node 5) is non-null, all of its ancestors are non-null. The other nodes in the tree are null.
Notice that the probability of making a false positive error in this tree depends on whether the hypothesis of no effect on a null node like node 3 or 13 is (1) tested and (2) rejected (we are using "null node" as a short hand for "node with no causal effect" or "node where the hypothesis of no effects is true"). Figure 4 contains a tree where we want to reject the root node. In fact, we should reject that node with probability much greater than $\alpha$ if we have enough data and the causal effect aggregated to that node is large.
This line of reasoning reveals that the probability of false positive errors depends on the power of the test at the root node (we write the statistical power of a test at node $i$ as $\pi_{i}$). Consider two extreme cases: (1) the overall test has very low power and (2) it has very high power. If the overall test has very low power, say $\pi_{1}\leq\alpha$, then this test will reject very rarely, and, in turn, the tree will show very few false positive errors. In fact, if $\pi_{i}\leq\alpha$, then the two conditions used above control the FWER in the strong sense: not rejecting at the root controls the FWER whether the lack of rejection is because all blocks have no effects and the test has false positive error rate of less than $\alpha$ or because at least one block has a non-zero effect but the test lacks the power to detect that effect ($\pi_{1}<\alpha$).
What about when $\pi_{1}>\alpha$? In this case, the multiple testing of null nodes are no longer protected by the root. For example, in Figure 4 rejection of the root will lead to 3 tests, two of them of null nodes. We will have two chances to falsely reject the null of no effects and, if the tree stopped with nodes 2,3 and 4, the FWER would be greater than $\alpha$ without further assumptions.
3.1 Strong control via error load and adaptive adjustment
Our approach exploits the tree’s gating structure. We show in Proposition 1 in the Supplement that the probability of a false positive error at any node is determined by the probability that the procedure both reaches that node and then rejects it.
Let $\mathcal{H}_{0}$ denote the set of indices $i$ for which $H_{i}$ is true, and define the conditional rejection probability $\theta_{j}$ at node $j$ as $\alpha_{j}$ (the test size) when $H_{j}$ is true, or $\pi_{j}$ (the power) when $H_{j}$ is false — in both cases conditional on all ancestors of $j$ having been rejected. Write $\text{anc}(i)$ for the set of proper ancestors of node $i$ (excluding $i$ itself). And let $R_{i}=\mathbb{I}\{\text{reject }H_{i}\}$: the rejection indicator at node $i$. Under the Stopping Rule, $R_{i}=1$ requires that $R_{j}=1$ for every $j\in\text{anc}(i)$. Then the FWER satisfies
$\mathrm{FWER}=\mathbb{P}\left(\bigcup_{i\in\mathcal{H}_{0}}\{R_{i}=1\}\right)\leq\sum_{i\in\mathcal{H}_{0}}\alpha_{i}\prod_{j\in\mathrm{anc}(i)}\theta_{j},$ (1)
and we define the product over an empty set of ancestors (for the root node) to be $1$.
Corollary 2 shows how this expression recovers weak FWER control when the root has very low power ($\pi_{1}\leq\alpha$): the root itself rarely rejects, so no descendant is tested. The expression also reveals the problem when the root has high power — the scenario we prefer. When $\pi_{1}=1$, rejection at the root is certain, and the FWER at level 2 is approximately $k\cdot\alpha_{2}$, the sum of the false positive rates across all $k$ children in that first branch of the tree from the root. With 10 children each tested at $\alpha=0.05$, the FWER reaches $0.5$.
To build intuition, imagine that the root has perfect power ($\pi_{1}=1$) and that the blocks are split into $k$ equal-sized groups at each level. Because each group contains $1/k$ of the data, the test at each child node has lower power than its parent: if we approximate power using the Normal distribution, $\theta\approx\Phi(\hat{\delta}\sqrt{N_{\text{node}}/4}-z_{\alpha/2})$, then halving the sample size at each split reduces power geometrically. At level $\ell$, the tree has $k^{\ell-1}$ nodes, and each null node is reached only if every ancestor on its path rejected. The product of these rejection probabilities along the path from the root determines how likely the procedure is to reach — and therefore to falsely reject — each null node.
We formalize this as the error load at level $\ell$: $G_{\ell}=k^{\ell-1}\prod_{j=0}^{\ell-1}\theta_{j}$, which is the product of the number of nodes at level $\ell$ and the probability that the procedure reaches that level (see Appendix B, Supplement). The error load measures how many null nodes at each level are exposed to testing. When the total error load $\sum_{\ell=1}^{L}G_{\ell}\leq 1$, the FWER contribution from all levels combined stays below $\alpha$, and the unadjusted procedure controls FWER (Theorem 2). We call this natural gating: power decays as data splitting reduces sample sizes at each level, and this decay limits how many null nodes the procedure can reach. In the multi-site education policy trials that motivate this paper — moderate effect sizes, $k\approx 3$, $L\approx 3$ — the error load is typically below 1 and no adjustment is needed.
When the error load exceeds 1 — as it will with wide trees, shallow depth, or high root power — we develop an adaptive $\alpha$-adjustment in Appendix B.4 (Supplement). If the test statistic has an asymptotically Normal distribution — Appendix C (Supplement) introduces one that does, following [undefl] — then we can compute power at each depth before testing begins and adjust $\alpha$ accordingly (Remark 8 and Theorem 9) as follows:
$\alpha_{\ell}^{\text{adj}}=\min\left\{\alpha,\frac{\alpha}{k^{\ell-1}\cdot\prod_{j=1}^{\ell-1}\hat{\theta}_{j}}\right\}$ (2)
We derive this expression for regular $k$-ary trees where each node has $k$ children. For irregular trees — where the number of children varies across nodes — the error load at each depth becomes a sum over the actual nodes at that depth, $G_{\ell}=\sum_{i\in\text{level }\ell}\prod_{j\in\text{anc}(i)}\hat{\theta}_{j}$, and the adjusted $\alpha$ at depth $\ell$ is $\min\{\alpha,\,\alpha/G_{\ell}\}$. The manytestsr R package implements this general formula via alpha_adaptive_tree(), which accepts a table of node sizes and computes per-depth power estimates for arbitrary tree structures.
The practical implication is that adjustment should be stringent near the root, where power is high and many null siblings are exposed, and relaxed at deeper levels, where power decay already gates the procedure. This resembles alpha-investing [undefb], but here the investment schedule is determined by the relationship between power and tree structure rather than by observed $p$-values.
3.1.0.1 Branch pruning.
The adaptive adjustment of Equation (2) computes the error load from the full tree before any testing begins. But when a branch fails to reject at depth $\ell$, the entire subtree below it is pruned: those nodes will never be tested, and the null sibling groups they would have contributed to deeper levels vanish from the error load. Branch pruning exploits this by recomputing the adaptive $\alpha$-schedule on the surviving subtree after each depth. When one of $k$ branches at depth 1 survives, the error load at depth 2 drops by a factor of roughly $k$, and the adjusted $\alpha$ at deeper levels increases correspondingly. The gain is largest in wide trees where most branches are null — exactly the setting where the pre-computed adjustment is most conservative. [^6] Note to the reader: We are still working on the proof for branch pruning. We present the simulations below which show it working, but this is a late breaking advance.
3.2 An Example: A Simulated Version of The Detroit Promise Program
The Detroit Promise Program (DPP) [undefr] involved enhanced advising, financial support, and encouragement of full-time enrollment to community college students in an effort to increase graduation rates. The trial randomized this bundle of treatments within five community colleges in the Detroit area: Henry Ford Community College, Macomb Community College, Oakland Community College, Schoolcraft College, and Wayne County Community College District. Each college fielded the trial for 3 academic year cohorts, and within the cohorts students were further subdivided into 1 to 4 experimental blocks within which the intervention was randomly assigned. To illustrate the algorithm, we generated a design that preserved the DPP structure — 5 college nodes under the overall node with varying numbers of cohorts and sub-cohort blocks, for 44 blocks total — but differed in ways that facilitated simulation. Every block had 50 students, and potential outcomes under control were drawn from $N(10,9)$ to roughly match the observed distribution of credits taken, but with more variation. We concentrated all non-zero effects in a single college: 9 blocks within HFCC have treatment effects while the 35 blocks across the other four colleges are pure null. This concentration mimics a pattern common in multi-site trials and visible in the MDRC data (Section 4): effects that are localized within particular sites rather than uniformly spread. Each block contains 50 students. For the illustration we use a large effect (Cohen’s $d=0.80$) within each of the 9 blocks with non-null effects so that the root-level test rejects and the tree structure is visible; the simulation study that follows evaluates realistic effect sizes.
Figure 5 shows the results of the structured testing approach with nodes that are truly non-zero displayed in blue. Any ancestor of a non-zero block is itself non-zero — a rejection at the college or cohort level is a true discovery, not an error. The algorithm tested 18 of the 44 blocks rather than all 44, pruning branches rooted at colleges where the college-level test did not reject. It correctly identified HFCC as the source of treatment effects and descended into its cohorts and blocks, detecting four of the nine non-null blocks while falsely rejecting two of the 35 null blocks.
Figure 5: Results of top-down testing in a simulation of 44 experimental blocks following the pre-specified experimental design of the Detroit Promise Program [undefr]. Nine blocks within HFCC have non-zero effects (Cohen’s $d=0.80$); all other blocks are pure null. The algorithm identifies HFCC and descends into its cohorts and blocks while pruning null colleges. Abbreviations: Henry Ford Community College (HFCC), Macomb Community College (MCC), Oakland Community College (OCC), Schoolcraft College (SC), Wayne County Community College District (WCC). Blue nodes have non-zero causal effects.
This illustration shows that a policy maker seeking to identify which colleges or blocks showed treatment effects would find the bottom-up approach disappointing: the [undefj] adjustment across 44 blocks detects just one block despite nine having genuine effects. The structured approach leads attention away from three of the five colleges and identifies not only more individual blocks but also higher-level groupings — the HFCC college itself — that can direct further exploration toward understanding why effects concentrated there [undefk, see].
The figure also shows the cost of using no $\alpha$-adjustment: Schoolcraft College (SC), a purely null college, rejected at the college level ($p=0.020$), and the procedure descended through SC Cohort 1 ($p=0.001$) to falsely reject two null blocks (Block0108 and Block0109, both $p=0.016$). These false rejections illustrate the cascade that the adaptive $\alpha$-adjustment of Equation (2) is designed to prevent. With adjustment, the SC college-level test would face a tighter threshold and would not reject, eliminating the entire false-rejection chain.
This is a single simulated dataset with a large effect size chosen to make the tree structure visible. Power and error rates describe the behavior of a procedure across repetitions of design. We now turn to a simulation at realistic effect sizes but using regular $k$-ary trees for simplicity to evaluate these properties systematically.
3.3 Does the approach control the FWER?
We used the same DPP design — 9 non-null blocks concentrated in one college, 35 pure-null blocks across the other four colleges — but set the effect size near the median of the observed overall MDRC effect sizes (Section 4) at Cohen’s $d=0.20$ for each of the 9 non-null blocks. So, this is a simulation with more power at the root than we would see in those studies. We repeated the simulation 10,000 times, re-randomizing treatment within each block while maintaining the same effect structure. Each iteration recorded whether a false positive error occurred at the level of individual blocks (“leaves”) or at any node in the tree. Table 2 summarizes the results.
Test Characteristic 3 Rules + Loc. Hom. + Loc. BH + Adapt. $\alpha$ + Ad. $\alpha$ Pr.
Nodes tested 3.983 3.350 3.484 3.077 3.521
Node FWER 0.028 0.005 0.005 0.005 0.009
Node False Rej. Prop. 0.018 0.003 0.003 0.003 0.006
Node Power 0.039 0.016 0.018 0.018 0.033
Leaves tested 2.447 1.989 2.098 1.776 2.120
Leaf Power 0.039 0.016 0.018 0.018 0.033
Leaf True Rejections 0.274 0.108 0.121 0.116 0.240
Leaf FWER 0.028 0.005 0.005 0.005 0.009
Leaf False Rej. Prop. 0.004 0.001 0.001 0.001 0.001
Bottom Up Power 0.004 0.004 0.004 0.004 0.004
Bottom Up True Rejections 0.033 0.036 0.038 0.037 0.035
Bottom Up FWER 0.024 0.022 0.028 0.022 0.023
Bottom Up False Rej. Prop. 0.001 0.001 0.001 0.001 0.001
Table 2: Operating characteristics of five testing procedures across 10,000 simulations ($\alpha=.05$) using simulated Detroit Promise Program data where 9 of 44 blocks (all in one college) have non-zero effects (Cohen’s $d=0.20$). ‘Node’ refers to any test in the tree (internal or leaf); ‘Leaf’ refers to individual experimental blocks; ‘Bottom Up’ shows the flat Hommel adjustment across all 44 blocks. ‘FWER’ is the proportion of simulations with at least one false rejection. ‘False Rej. Prop.’ is the mean proportion of truly null hypotheses falsely rejected per simulation. ‘Power’ is the mean proportion of non-null hypotheses correctly rejected. ‘True Rejections’ is the mean number of non-null hypotheses rejected. Boldface marks values discussed in the text.
Table 2 compares five top-down variants against a bottom-up Hommel baseline. The columns represent different combinations of the tools introduced above: “3 Rules” applies only the stopping rule and valid-test conditions (Conditions 1 and 2); “+ Loc. Hom.” and “+ Loc. BH” additionally adjust $p$-values within each sibling group using the [undefj] or [undef] procedure before comparison to $\alpha$; “+ Adapt. $\alpha$” uses the adaptive $\alpha$-adjustment of Equation (2); and “+ Ad. $\alpha$ Pr.” combines the adaptive $\alpha$-adjustment with branch pruning — recomputing the $\alpha$-schedule on the surviving subtree after each depth, so that dead branches free up alpha for the remaining tests. Each row reports a different operating characteristic: FWER (the proportion of simulations with at least one false rejection), power (the proportion of non-null hypotheses correctly rejected), and true rejections (the average number of non-null hypotheses rejected per simulation). The bottom three rows show the corresponding quantities for the flat Hommel adjustment applied to all 44 blocks simultaneously.
At $d=0.20$, the error load is 4.74 — well above the threshold of 1 where Theorem 2 guarantees control. Yet the unadjusted two-conditions procedure still controls the FWER at 0.031, because the error-load condition is sufficient but not necessary: the DPP tree’s concentrated-effect structure provides additional natural gating beyond what the general bound captures. The adaptive $\alpha$-adjustment (Equation 2) brings the FWER down further to 0.006. Bottom-up Hommel also controls the FWER (0.022) but at a steep cost: it detects effects in only 0.4% of the truly non-null blocks, compared to 3.9% for the unadjusted top-down procedure and 2.0% for the adaptive variant.
Although absolute power is low — realistic at $d=0.20$ with $N=50$ per block — the relative advantage is large. The unadjusted structured approach averages 0.27 true leaf rejections per simulation, nearly eight times the bottom-up Hommel’s 0.035. Even the most conservative structured variant (adaptive $\alpha$) averages 0.13 true rejections, more than three times the bottom-up rate. Branch pruning improves further: the pruned adaptive variant averages 0.24 true rejections — nearly twice the non-pruned adaptive rate — because once the four null colleges fail to reject at depth 1, their subtrees vanish from the error load and the $\alpha$-schedule relaxes for deeper levels within HFCC. At $d=0.30$, pruning yields 0.61 true rejections versus 0.21 for the non-pruned adjustment, a gain of $2.9$ times.
This advantage arises because the top-down procedure aggregates signal at the college level ($N=450$ for HFCC) before descending to individual blocks, while the bottom-up approach tests each block with less information — $N=50$ — and pays a heavy multiplicity penalty across all 44. Branch pruning amplifies this advantage: when the procedure correctly identifies HFCC as the active college and prunes the four null colleges, the effective tree narrows from 5 branches to 1 and the error load drops accordingly.
The table also shows two local-adjustment variants. The “plus Local Hommel” column applies a Hommel adjustment to the children of each internal node: leaf power drops from 0.040 to 0.015 and the FWER from 0.031 to 0.004. The “plus Local BH” column applies the Benjamini-Hochberg procedure locally: similar power (0.017) and FWER (0.005). These local adjustments provide additional conservatism, but at $d=0.20$ the unadjusted procedure already controls the FWER, and the pruned adaptive $\alpha$ achieves comparable FWER control (0.009) with the highest power among the adjusted variants (0.034).
The DPP results illustrate a pattern that may hold broadly: the error-load condition of Theorem 2 is conservative, and the unadjusted procedure often controls the FWER even when the error load exceeds 1. But how far can this tolerance stretch? At $d=0.30$ with the same DPP tree, the error load reaches 8.85 and the unadjusted FWER rises to 0.039 — still below 0.05, but close. The adaptive $\alpha$-adjustment provides insurance, holding the FWER at 0.003 with a power cost (leaf power drops from 0.098 to 0.031). Branch pruning recovers much of this lost power: the pruned variant achieves a FWER of 0.009 with leaf power of 0.079. In wider trees with stronger effects, the unadjusted procedure will eventually inflate beyond 0.05 and $\alpha$-adjustment is necessary.
We now probe this boundary with a simulation study that varies tree width, effect size, and the proportion of null hypotheses to create scenarios spanning error loads from 0.2 to 3.1. This design isolates the testing procedure from data-generation details, letting us identify where the unadjusted procedure breaks down and how much power the adaptive correction costs.
3.3.1 Simulation Study of Strong FWER Control
The weak-control simulation above tested trees under the global null, where every hypothesis is true. We now study how the procedure behaves when some hypotheses are false — the setting in which strong FWER control matters. The central question is whether the unadjusted top-down procedure inflates the FWER when powerful root tests expose many null sibling groups to testing, and whether the adaptive $\alpha$-adjustment of Equation (2) corrects the inflation.
We consider two tree structures: binary ($k=2$, depth 3, yielding 8 leaves) and quaternary ($k=4$, depth 3, yielding 64 leaves). For each tree we vary the effect size and the proportion of null hypotheses, creating 14 scenarios in total. We express effect sizes as Cohen’s $d$ — the same planning parameter used to calibrate the adaptive $\alpha$-adjustment. Three values calibrate three power regimes: $d=0.04$ yields root power of approximately 0.15 (error load well below 1, natural gating suffices); $d=0.10$ yields root power of approximately 0.61 (error loads near the transition threshold); and $d=0.15$ yields root power of approximately 0.92 (error load above 3, the regime where we expect the unadjusted procedure to inflate). Each effect size is crossed with two null proportions: 80% null (the worst case for inflation, since many null sibling groups are exposed) and 50% null. Two additional all-null scenarios ($k=2$ and $k=4$) verify that all methods control the FWER when every hypothesis is true.
For each scenario, we draw $p$-values rather than simulating raw data, isolating the testing procedure from data-generation details. Non-null leaves receive $p$-values drawn from $\text{Beta}(a,1)$, where the shape parameter $a$ is calibrated so that the rejection probability matches the power implied by Cohen’s $d$ and the sample size available at that leaf. Null leaves receive $p$-values from $U(0,1)$. Parent and child $p$-values are drawn independently given each node’s null status; the procedure does not require monotonicity between parent and child $p$-values (see Remark 9 in the Supplement).
We compare seven methods organized into two families:
Top-down, unadjusted (TD-Unadj).
The basic gating procedure of Section 2, testing each node at the nominal $\alpha=0.05$ with no local multiplicity adjustment.
Top-down, Hommel (TD-Hommel).
Same gating rule, but within each sibling group the $p$-values are adjusted by the Hommel procedure before comparison to $\alpha$.
Top-down, adaptive (TD-Adapt).
The gating procedure with $\alpha$-levels adjusted by Equation (2), using estimated power at each depth. No local adjustment.
Top-down, adaptive + Hommel (TD-Adapt-Hom).
Both adaptive $\alpha$-adjustment and local Hommel correction.
Top-down, adaptive pruned (TD-Adapt-Pr).
The adaptive $\alpha$-adjustment with branch pruning: after each depth, the error load is recomputed on the surviving subtree and the $\alpha$-schedule for deeper levels is updated.
Bottom-up, Hommel (BU-Hommel).
All leaves are tested simultaneously, with the Hommel procedure applied globally.
Bottom-up, BH (BU-BH).
All leaves are tested simultaneously, with the Benjamini–Hochberg procedure applied globally.
For each of the 14 scenarios and 7 methods, we repeat the procedure 10,000 times. Within each scenario, all five top-down methods use the same random seed, so they operate on identical trees and leaf $p$-values. We record the FWER (the proportion of simulations in which at least one true null hypothesis is rejected) and power (the proportion of non-null leaves correctly rejected, averaged across simulations).
Table 3 reports the FWER for all seven methods across the 14 scenarios. Table 4 compares the number of true discoveries between the adaptive top-down procedures (with and without branch pruning) and the bottom-up Hommel alternative.
FWER
$k$ $d$ Null $\sum G$ TD TD-Hom TD-Adp TD-A-H TD-A-Pr BU-Hom BU-BH
2 0.04 50% 0.2 0.000 0.000 0.000 0.000 0.000 0.024 0.026
2 0.10 50% 1.2 0.009 0.004 0.011 0.003 0.014 0.026 0.031
2 0.15 50% 3.1 0.042 0.020 0.016 0.007 0.037 0.024 0.034
2 0.04 80% 0.2 0.005 0.002 0.005 0.001 0.007 0.044 0.045
2 0.10 80% 1.2 0.034 0.018 0.028 0.013 0.033 0.042 0.044
2 0.15 80% 3.1 0.075 0.056 0.035 0.023 0.053 0.045 0.050
2 100% 1.2 0.050 0.051 0.052 0.052 0.051 0.051 0.052
4 0.04 50% 0.2 0.001 0.000 0.000 0.000 0.001 0.027 0.028
4 0.10 50% 1.3 0.014 0.001 0.004 0.000 0.011 0.021 0.022
4 0.15 50% 3.1 0.057 0.003 0.004 0.000 0.026 0.024 0.026
4 0.04 80% 0.2 0.001 0.000 0.001 0.000 0.003 0.040 0.042
4 0.10 80% 1.3 0.023 0.001 0.008 0.000 0.030 0.041 0.043
4 0.15 80% 3.1 0.115 0.009 0.010 0.001 0.054 0.041 0.043
4 100% 1.3 0.048 0.050 0.049 0.052 0.047 0.047 0.048
Table 3: Strong FWER control across 14 scenarios and 7 methods (10,000 simulations per cell; simulation error $\approx 0.010$). TD = top-down with nominal $\alpha=0.05$; Hom = local Hommel adjustment within each sibling group; Adp = adaptive $\alpha$-adjustment (Equation 2); A-H = both adaptive and Hommel; A-Pr = adaptive with branch pruning; BU-Hom/BH = bottom-up Hommel/Benjamini–Hochberg applied to all leaves. Boldface indicates FWER exceeding $\alpha$ plus simulation error. $d$ = Cohen’s $d$ (effect size); Null = proportion of null hypotheses; $\sum G$ = total error load. All-null rows (100% null) use a placeholder effect size.
TD-Adp TD-A-Pr BU-Hom
$k$ $d$ Null $\sum G$ FWER Disc. FWER Disc. FWER Disc. Ratio
[…]
Table 4: True discoveries and testing efficiency: top-down adaptive (TD-Adp), top-down adaptive with branch pruning (TD-A-Pr), and bottom-up Hommel (BU-Hom) across 12 scenarios with non-null effects (10,000 simulations each). TD methods count correctly rejected non-null hypotheses at all levels of the tree; BU-Hom counts correctly rejected non-null leaves. Disc. = average true discoveries per simulation. Ratio = TD-A-Pr discoveries $\div$ BU-Hom discoveries. Boldface FWER indicates inflation beyond $\alpha+$ simulation error.
The rows labeled “—” in the $d$ column of Table 3 are the all-null scenarios ($k=2$ and $k=4$ with 100% null hypotheses). These rows serve as a sanity check: they reproduce the weak-control setting of Table 1 within the strong-control simulation framework. All methods control the FWER at the nominal level in these rows, as expected.
Three patterns emerge from these tables. First, when the total error load $\sum G_{\ell}$ is at most 1 — the regime where Theorem 2 guarantees control — every method controls the FWER at or below the nominal level. The unadjusted top-down procedure, the adaptive variant, and both bottom-up procedures all remain below 0.05. This confirms that natural gating works in the regime for which it was designed. Recall that the error load measures how many null nodes at each depth are exposed to testing: it is the product of the number of nodes at that depth and the probability that the procedure reaches them (a function of power along the path from the root). When the error load is small, power decay from data splitting prevents the procedure from reaching enough null nodes to inflate the FWER.
Second, when the error load rises above 1, the unadjusted top-down procedure begins to inflate. At $k=4$, $d=0.15$, and 80% null hypotheses — a total error load of 3.1 — the unadjusted FWER reaches 0.115, more than double the nominal level. The adaptive $\alpha$-adjustment reduces this to 0.010 in the same scenario; branch pruning brings it to 0.054. The local Hommel adjustment alone brings the FWER to 0.009, and combining adaptive adjustment with Hommel yields 0.001. The adaptive adjustment works because it tightens the threshold at shallow depths where root power exposes many null sibling groups, then relaxes it at deeper levels where power decay already gates the procedure. Branch pruning is less conservative than the non-pruned adjustment but more aggressive, recovering alpha from dead branches to test deeper levels more liberally.
[…]
The practical recommendation is straightforward: compute the error load before testing. If $\sum G_{\ell}\leq 1$, the two conditions of Section 2 suffice. If $\sum G_{\ell}>1$, apply the adaptive $\alpha$-adjustment with branch pruning. In either case, the top-down procedure provides substantially more power than a bottom-up correction while controlling FWER at the same rate.
4 Application: The MDRC RCT Data
From 2003 to 2019 the MDRC organization worked with community colleges to field randomized field trials assessing the effects of interventions like financial support or improved advising meant to improve the educational outcomes of community college students. MDRC deposited data from 31 of these studies in the ICPSR data repository, where they are available under restricted data use agreements [undefa]. We applied the method to the 25 studies that randomly assigned an intervention to individual students within strata, following each study’s pre-specified series of splits. The outcome in each study is the number of community college credits taken in the first main session after the intervention — for example, credits taken during the next full academic year [undefac, undefv].
Table 5 reports the 12 studies where at least one overall test (Wilcoxon rank or t-test) rejected at $\alpha=0.05$; the remaining 13 studies appear in Supplement Table D.3.
[…]
Table 5: Structured (top-down) vs. bottom-up testing in the 12 MDRC studies where at least one overall test rejected at $\alpha=0.05$. ’3 Rules’ applies the stopping rule and valid tests only (weak FWER control). ’+ Adj $\alpha$ Pr.’ adds the adaptive $\alpha$-adjustment with branch pruning (strong FWER control). ’Ratio’ is the number of top-down node detections divided by the number of bottom-up Hommel detections ($\infty$ when Hommel finds nothing but the tree does; ’—’ when neither finds anything). The remaining 13 studies (overall $p>0.05$) are reported in Supplement Table D.3.
We used the Wilcoxon rank test for all tests and show the overall t-test only as a comparison with published results; these data are very skewed and the rank test generally has higher power in this setting. The distribution of the outcome variable (credits earned) is reported in Supplement Table D.2: the outcome is skewed toward 0, but with a handful of students taking many credits.
[…]
Across these 25 studies, the structured approach is at least as useful as existing methods and usually more powerful, with at least the same error rate control as the FDR-oriented BH approach. The DPP simulation above suggests that strong FWER control holds in designs of this size.
5 Limitations
At least three open questions deserve attention.
5.0.0.1 Test statistic.
The structured testing approach requires a test at each node that can detect any departure from the null, not just a mean shift. If half of the blocks have a large negative effect and half a large positive effect, a difference-in-means test at the root could fail to reject. [^8] Thanks to Kevin Quinn for this question. We address this by creating a test statistic based on energy statistics [undefs, undefy] combined with the multivariate permutation framework of [undefl] and related to the $d^{2}$-test of [undefg]. This statistic is sensitive to distributional differences of any kind and is described in Appendix C (Supplement) and implemented in the manytestsr R package. [^9] The manytestsr package is available at https://bowers-illinois-edu.github.io/manytestsr/. The combined Stephenson rank statistic of [undefm] is a promising alternative that merits investigation. [^10] The R package for that statistic is at https://bowers-illinois-edu.github.io/CMRSS.
5.0.0.2 Data-dependent splitting.
The examples in this paper use tree structures pre-specified by the experimental design. Many block-randomized experiments lack such administrative hierarchies. We have implemented several data-based splitting algorithms in the manytestsr package — for example, k-means clustering of blocks by covariate distributions or splits that equalize the number of units per group. Preliminary simulations show that these algorithms control the FWER and provide more power than bottom-up approaches, but their systematic evaluation is future work.
5.0.0.3 FDR control.
This paper focuses on FWER control. An alternative is to allocate the error budget dynamically as tests descend the tree, targeting control of the false discovery rate (FDR) rather than the FWER. The alpha-investing procedure of [undefb] and its refinements [undefq, undefaa] provide a natural framework for this extension. FDR-oriented tree testing would offer different guarantees and different power profiles and deserves its own treatment.
6 Discussion
An analyst with a block-randomized experiment and a pre-specified hierarchy can now do the following: test the overall null at the root, divide the blocks into groups, and continue testing down the tree, stopping in any branch where the null is not rejected. If only weak FWER control is needed — the standard in exploratory research — the stopping rule and valid-test conditions suffice, and no further adjustment is required. If strong FWER control is needed, one checks the error load: when power decays fast enough relative to the branching factor, the unadjusted procedure already controls FWER in the strong sense (Theorem 2). When the error load exceeds 1, the adaptive $\alpha$-adjustment with branch pruning compensates for the many simultaneously tested families while recovering alpha from branches that fail to reject (Remark 8 and Theorem 9). The manytestsr R package implements these approaches.
The method works best when the experimental design provides a natural hierarchy and blocks are numerous. With few blocks, the most powerful bottom-up adjustments like the Hommel procedure are nearly as powerful, because the tree is shallow and the structured approach has little room to exploit the stopping rule. The gains are largest in experiments with dozens to hundreds of blocks organized in several levels — the setting of multi-site education trials, multi-center clinical trials, and large-scale field experiments. The current development of adaptive $\alpha$-adjustment requires estimates of power at each level, which depend on assumptions about effect sizes and the test statistic’s distribution; however, in many practical settings the error load is small enough that no adjustment is needed at all. When adjustment is needed, branch pruning provides a meaningful power gain — nearly doubling the number of true discoveries in the DPP simulation — at no cost to FWER control.
Several extensions remain. Data-dependent splitting algorithms would extend the method to experiments without administrative hierarchies and could connect this testing-based approach to other approaches aiming to understand how treatment effects vary by covariate values (§ 5). FDR-oriented variants using alpha-investing could offer more detections when researchers are willing to tolerate a higher false discovery rate.
The policy-maker who began by asking where effects occurred now has a principled answer. The experiment’s own structure — its sites, cohorts, and blocks — provides the hierarchy for testing, and the stopping rule provides the error control. What the standard adjustment discards as noise, the structured approach recovers as signal.
Appendix A Weak FWER Control
Intuition for the proof of Theorem 1 .
If all null hypotheses on the tree were true, such that no effects existed for any unit in any block, any rejection anywhere in the tree would be a false positive. The stopping rule states that we must first reject the overall null hypothesis of no effects at the root before we are allowed to test any other null hypothesis on the tree. So the event “at least one false rejection occurs” is contained in the event “the root is rejected.” If the root test at level $\alpha$ has size $\leq\alpha$, then the probability of doing any other test on the tree is at most $\alpha$. Hence the weak FWER is $\leq\alpha$.
A.1 Conditions on the Procedure
Condition 3 (Stopping rule).
Test hypothesis $H_{i}$ only if every ancestor of node $i$ has been rejected. If the root hypothesis is not rejected, no further tests are performed. If any hypothesis on a path from root to leaf is not rejected, no descendant on that branch is tested at all. In Figure 2, we do not test $H_{5}$ unless we have rejected both $H_{2}$ and $H_{1}$.
This condition distinguishes our procedure from tree-structured testing frameworks that test every node and then determine which rejections to keep [undefae, undefd]. Our procedure may test only a small fraction of the tree’s nodes — and each untested node is a test that cannot produce a false positive.
Condition 4 (Valid tests).
Each test at a node, if executed on its own, has a false positive rate no greater than $\alpha$. For proof of weak control of the FWER this condition implies that the test at the root is level $\alpha$, i.e., its size is $\leq\alpha$ when the root null is true.
Remark 1.
In block-randomized experiments, randomization-based tests satisfy the Valid Tests Rule by design: permuting the treatment assignment within blocks produces $p$-values with guaranteed size control [undefu, § 2.4]. All tests that we consider in this paper are randomization-based tests justified by the design of the randomized experiments to which they are applied.
{theorem}
[Conditions 1 and 2 suffice for weak FWER control; restated from main text]
A family of true null hypotheses organized on a regular or irregular $k$-ary tree and tested following the stopping rule (Condition 1) with valid tests at each node (Condition 2) will produce a family-wise error rate no greater than $\alpha$. We call a $k$-ary tree “regular” if the number of child nodes of a parent is the same for all nodes in the tree. A $k$-ary tree is “irregular” if the number of child nodes of a parent node may differ within and across levels of the tree.
A.2 Proof: Weak Control of the FWER Using Two Conditions
Proof of Theorem 2.
We assume that all hypotheses on the tree are true. We write $V$ for the number of false positive rejections arising from the testing procedure on the whole tree, $E$ as the event that the hypothesis on the root node of the whole tree is rejected and $E^{c}$ as the event that the hypothesis on the root node is not rejected (the complement of $E$).
We need to check
$\mathbb{P}(V\geq 1)\leq\alpha,$ (A.2)
under Conditions 1 and 2.
Note that we can decompose the probability of at least one false positive rejection into two parts depending on whether the root-node test was rejected, $E$, or not, $E^{c}$:
$\mathbb{P}(V\geq 1)$ $=\mathbb{P}(V\geq 1|E)\mathbb{P}(E)+\mathbb{P}(V\geq 1|E^{c})\mathbb{P}(E^{c}),$
and since we cannot have a false rejection without a rejection, the second term is 0, so
$=\mathbb{P}(V\geq 1|E)\mathbb{P}(E).$
Note that $\mathbb{P}(V\geq 1|E^{c})=0$ by our Stopping Rule: a failure to reject at the root node means (1) no false positive error at the root node (so $V=0$) and (2) that no other hypotheses are tested and thus no further rejections are possible and thus no false rejections are possible in this case.
It remains to bound the right-hand side. Since all null hypotheses are true, $\mathbb{P}(E)\leq\alpha$ by Valid Tests (the probability of rejecting a true null is at most the level of the test). And $\mathbb{P}(V\geq 1\mid E)\leq 1$ because it is a probability. Therefore,
$\mathbb{P}(V\geq 1)=\mathbb{P}(V\geq 1|E)\mathbb{P}(E)\leq\mathbb{P}(E)\leq\alpha,$
which completes the proof of (A.2). ∎
Remark 2.
Note that this proof does not require the sample splitting step used in the algorithm of the paper. Weak control can be shown to hold for hypotheses organized into a tree-like structure even without the sample reduction at each level of the tree.
Remark 3 (Relationship to prior work).
The conclusion that a stopping rule combined with valid individual tests suffices for weak FWER control is not surprising given the existing literature, but the specific proof above—a direct application of the law of total probability, conditioning on the root—does not appear in prior work in this form.
[undeft] establishes the analogous result for linear sequences of hypotheses: testing in a fixed order and stopping at the first non-rejection controls the FWER at $\alpha$. His argument is essentially the same gating logic, but applies only to chains, not trees with branching.
[undefp] proves FWER control for tree-structured hypotheses via a different route entirely: a Bonferroni bound over disjoint maximal true-null clusters. That proof exploits the hierarchy to partition the null hypotheses, rather than conditioning on the root.
[undefe] prove a general sequential rejection principle: any procedure satisfying monotonicity and a single-step Bonferroni condition controls FWER. Our stopping-rule procedure is a special case of their framework, but their proof proceeds by induction over rejection sets and does not isolate the direct gating argument used here. Similarly, the inheritance procedure of [undefd] verifies the Goeman–Solari conditions for a specific tree algorithm, but again uses the general-purpose machinery rather than the one-step decomposition.
In short, the result is a known consequence of principles in the literature. The proof is our own formulation, chosen for transparency: it makes visible exactly where the two conditions—stopping rule and valid tests—do their work, with no additional apparatus (no sample splitting, no monotonicity, no abstract critical-value conditions).
Appendix B Strong FWER Control
Intuition
Weak FWER control (Theorem 2 in the main text) guarantees that the procedure’s false-rejection rate is at most $\alpha$ when every null hypothesis is true. Strong FWER control extends this guarantee to any configuration of true and false nulls — including settings where some nodes have genuine effects and others do not. Whether strong control holds, and how conservative a procedure must be to achieve it, depends on how much of the tree the procedure actually explores. And exploration of the tree depends on the power of the tests as it changes throughout the tree as the overall sample is reduced via splitting.
For example, if one experimental block out of 100 has a true non-null causal effect, then we would hope to reject the test of the null of no effects at the root node — only a rejection at the root would allow us any hope of descending to test and reject the null hypothesis of no effects within that one block. But, if the test has low power at the root (say, the sample size is small or the effect is weak and power $\leq 0.05$), then the procedure will not reject the null, no further tests will occur, and the FWER will be $\leq 0.05$. That is, the probability of making a false positive error at any node depends both on reaching that node and on falsely rejecting that node. And the FWER thus depends on the number of falsely rejected nodes. We here develop a procedure that adapts the rejection level $\alpha$ to the contours of the tree.
The argument proceeds in four steps:
  1. 1.
    We derive a general expression for the FWER as a sum of per-node contributions, each discounted by the probability that the procedure reaches that node (Proposition 1).
  2. 2.
    We group those contributions by tree depth and identify the structural quantity — the error load — that governs whether the FWER stays below $\alpha$ (Theorem 4 and Proposition 2).
  3. 3.
    We establish conditions under which power decay alone keeps the error load small enough that no adjustment is needed — a phenomenon we call natural gating (Section B.3).
  4. 4.
    For trees where natural gating fails, we develop the adaptive $\alpha$-adjustment and prove that it controls the FWER (Theorem 9 in Section B.4).
The reader primarily interested in the algorithm can skip to Section B.4; the intervening results build the intuition for why the adjustment takes the form it does.
B.1 Setup and Notation
The analysis in this section assumes both conditions on the testing procedure—the Stopping Rule (Condition 3) and Valid Tests (Condition 4)—stated in Appendix A. No additional conditions are needed: the proofs below rely on these two conditions together with the intersection structure of the tree (defined below).
Consider a rooted regular or irregular $k$-ary tree of depth $L$ (root at level 0, leaves at level $L$) with the following notation.
  • Each leaf of the tree corresponds to an independent experimental block (or finest unit of randomization). Each non-leaf node represents a group of blocks.
  • Each node $i$ has an associated null hypothesis $H_{i}$ about the units it contains. In this paper we consider only the null hypothesis of no effects: $H_{i}$ asserts that the treatment has no causal effect on any unit in the blocks represented by node $i$.
  • The hypotheses obey an intersection structure: each non-leaf hypothesis is the conjunction of its children. A parent hypothesis is true if and only if every one of its child hypotheses is true. This ensures that whenever a parent is false, at least one child must also be false. Concretely: if a single experimental block contains a genuine treatment effect, then every ancestor of that block also contains that effect, and the null hypothesis of no effects is false for them all. The false hypotheses therefore form a connected subtree rooted at the root.
  • $\mathcal{H}_{0}\subseteq\{1,\dots,N_{\text{nodes}}\}$: the index set of nodes whose null hypothesis $H_{i}$ is true (there are no effects in any block represented by node $i$).
  • $\text{anc}(i)$: the set of proper ancestors of node $i$ (excluding $i$ itself).
  • Testing proceeds top-down by the Stopping Rule (Condition 3): node $i$ is tested only if every ancestor of $i$ has already been rejected.
  • $R_{i}=\mathbb{I}\{\text{reject }H_{i}\}$: the rejection indicator at node $i$. Under the Stopping Rule, $R_{i}=1$ requires that $R_{j}=1$ for every $j\in\text{anc}(i)$.
  • $\alpha$: the nominal significance level (e.g. $0.05$). We require $\alpha_{i}\leq\alpha$ for every node $i$.
  • By Valid Tests (Condition 4), each node $i$ is tested by a randomization-based test with size $\alpha_{i}\leq\alpha$. That is, when $H_{i}$ is true,
    $\mathbb{P}(R_{i}=1\mid H_{i}\text{ true and all ancestors rejected})\leq\alpha_{i}.$
  • $\pi_{i}$: the conditional power at node $i$ when $H_{i}$ is false: $\pi_{i}=\mathbb{P}(R_{i}=1\mid H_{i}\text{ false},\ \text{all ancestors rejected})$. Power depends on the sample size available at node $i$ (which shrinks as we descend the tree) and on the magnitude of the treatment effect.
  • The conditional rejection probability $\theta_{j}$ combines the two cases above into a single quantity:
[…]
    In words: $\theta_{j}$ is the probability of rejecting $H_{j}$ given that the procedure has reached node $j$ (all ancestors rejected). When $H_{j}$ is false this probability is the power $\pi_{j}$; when $H_{j}$ is true it is the test size $\alpha_{j}$.
  • We call a true-null node exposed if its parent is non-null (the parent’s null hypothesis of no effects is false). Exposed nulls sit at the boundary between the false subtree and the true portion of the tree. They are the only true nulls that the top-down procedure can reach directly, because a true-null parent will almost certainly not be rejected, shielding all of its descendants from testing.
    We write $e_{\ell}$ for the number of exposed nulls at level $\ell$, and $m_{\ell}$ for the number of non-null nodes at level $\ell$.
B.2 FWER Decomposition
We derive an upper bound on the FWER in terms of the test sizes $\alpha_{i}$ and the conditional rejection probabilities $\theta_{j}$ along the tree.
Proposition 1 (FWER Expression for Sequential Tree Testing).
Consider the sequential tree testing procedure described above, and write $\mathcal{H}_{0}$ as the set of indices $i$ for which $H_{i}$ is true. Then the familywise error rate satisfies
$\mathrm{FWER}=\mathbb{P}\left(\bigcup_{i\in\mathcal{H}_{0}}\{R_{i}=1\}\right)\leq\sum_{i\in\mathcal{H}_{0}}\alpha_{i}\prod_{j\in\text{anc}(i)}\theta_{j},$ (B.4)
where $\theta_{j}$ is given by (B.3) and the product over an empty set of ancestors (for the root node) is $1$.
In words: each true null $i$ contributes to the FWER through two factors — the test size $\alpha_{i}$ (how likely the test is to falsely reject, given that it is actually run) and the path product $\prod_{j\in\text{anc}(i)}\theta_{j}$ (how likely the procedure is to reach node $i$ at all). Nodes buried deep in the tree, behind many ancestors with low rejection probability, contribute very little to the FWER — the procedure almost never reaches them.
The bound (B.4) is tightest when exposed true nulls are few and widely separated in the tree, and is exact when there is only a single exposed true null.
Proof.
Step 0: Define the FWER By definition,
$\mathrm{FWER}=\mathbb{P}\left(\bigcup_{i\in\mathcal{H}_{0}}\{R_{i}=1\}\right).$ (B.5)
Step 1: Decompose rejection at node $i$. By the Stopping Rule (Condition 3), we test $H_{i}$ only if all its ancestors have been rejected:
$\{R_{i}=1\}=\{R_{i}=1\}\cap\bigcap_{j\in\text{anc}(i)}\{R_{j}=1\}.$ (B.6)
Step 2: Bound the false-rejection probability at a single true null. We now focus on a single true null hypothesis $i\in\mathcal{H}_{0}$, because the FWER counts only false rejections — rejecting a node whose hypothesis is actually false would be a correct detection, not an error. For this true null $i$, we want to bound $\mathbb{P}(R_{i}=1)$, the probability that $i$ is falsely rejected.
The stopping rule means that $R_{i}=1$ can only happen if the procedure reaches node $i$, which requires all ancestors to be rejected first. Conditional probability lets us separate the act of reaching node $i$ (the ancestor product) from the act of falsely rejecting once there (the test size $\alpha_{i}$):
$\mathbb{P}(R_{i}=1)$ $=\mathbb{P}\left(R_{i}=1\,\middle|\,\bigcap_{j\in\text{anc}(i)}\{R_{j}=1\}\right)\mathbb{P}\left(\bigcap_{j\in\text{anc}(i)}\{R_{j}=1\}\right)$ (B.7)
$\leq\alpha_{i}\cdot\mathbb{P}\left(\bigcap_{j\in\text{anc}(i)}\{R_{j}=1\}\right),$ (B.8)
where the inequality uses Valid Tests (Condition 4): since $H_{i}$ is true, the conditional rejection probability is bounded by the size $\alpha_{i}$.
Step 3: Evaluate the probability that all ancestors are rejected. Step 2 left us with the ancestor-rejection probability $\mathbb{P}(\bigcap_{j\in\text{anc}(i)}\{R_{j}=1\})$ as an unexpanded term. We now evaluate it. Let the ancestors of node $i$ along the root-to-$i$ path be $a_{1},a_{2},\dots,a_{m}$, ordered from root downwards. By the chain rule of conditional probability, and because each $\theta_{j}$ is defined as the probability of rejecting at node $j$ given that all of $j$’s ancestors are already rejected, the product telescopes:
$\mathbb{P}\left(\bigcap_{j\in\text{anc}(i)}\{R_{j}=1\}\right)=\prod_{j\in\text{anc}(i)}\theta_{j}.$ (B.9)
Each factor $\theta_{j}$ in the product conditions on all prior rejections, so the chain rule applies directly. Substituting into (B.8):
$\mathbb{P}(R_{i}=1)\leq\alpha_{i}\prod_{j\in\text{anc}(i)}\theta_{j}\quad\text{for each }i\in\mathcal{H}_{0}.$ (B.10)
Step 4: Union bound. Steps 1–3 gave us a bound on $\mathbb{P}(R_{i}=1)$ for each individual true null $i$: the probability that the procedure reaches node $i$ and then falsely rejects it. The FWER asks about a different event — that at least one true null anywhere in the tree is falsely rejected. The union bound (also called Boole’s inequality) bridges this gap: the probability that at least one event in a collection occurs is at most the sum of their individual probabilities.
$\mathrm{FWER}=\mathbb{P}\left(\bigcup_{i\in\mathcal{H}_{0}}\{R_{i}=1\}\right)\leq\sum_{i\in\mathcal{H}_{0}}\mathbb{P}(R_{i}=1)\leq\sum_{i\in\mathcal{H}_{0}}\alpha_{i}\prod_{j\in\text{anc}(i)}\theta_{j},$ (B.11)
which is (B.4). The first inequality is the union bound; the second substitutes the per-node bound from (B.10). The union bound can overcount — if two sibling nulls are both falsely rejected in the same simulation run, that run is counted twice — but it never undercounts, so it gives a valid upper bound on the FWER. ∎
Remark 4 (When is the bound tight?).
The union bound is tightest when exposed true nulls are few and widely separated in the tree, and is exact when there is only a single exposed true null. It can be loose when multiple exposed true nulls cluster at the same level, because the union bound counts the possibility that two or more nulls are both falsely rejected in the same run, even though this event is very rare.
Example. Consider a tree with $k=3$ and $L=2$ where the root is false and two of its three children are true nulls, each tested at level $\alpha=0.05$. The union bound gives $\text{FWER}\leq 2\times 0.05=0.10$, while the exact probability is $1-(1-0.05)^{2}=0.0975$. The bound is nearly tight because simultaneous false rejection of both nulls is rare (probability $0.05^{2}=0.0025$). With more exposed nulls the gap widens: with $k=10$ and 9 exposed nulls, the union bound gives $9\times 0.05=0.45$ while the exact value is $1-(1-0.05)^{9}=0.37$. In both cases the bound is conservative (never too small), which is what a valid FWER guarantee requires.
The node-level bound in Proposition 1 expresses the FWER as a sum over individual nodes. To design a practical adjustment, we need a formula that depends on depth rather than on individual node identities — because the adaptive $\alpha$-schedule assigns one significance level per depth, not per node. Grouping terms in (B.4) by level reveals the structural quantities that govern the adjustment. The false nulls form a connected subtree rooted at the root (by the intersection structure), and if the root is true, the global null holds and weak FWER control (Theorem 1) already applies.
{theorem}
[FWER Decomposition by Level]
(a) The FWER contribution from the $e_{\ell}$ exposed nulls at level $\ell$ satisfies
$\text{FWER contribution from level }\ell\leq\left(e_{\ell}\cdot\alpha_{\ell}\cdot\prod_{j=1}^{\ell-1}\theta_{j}\right)$ (B.12)
where $\alpha_{\ell}$ is the common test size at level $\ell$, and the product $\prod_{j=1}^{\ell-1}\theta_{j}$ runs over the non-null ancestors along the path from the root to level $\ell$ — the probability that the procedure reaches level $\ell$.
(b) Writing $\mathcal{H}_{0}^{(\ell)}$ for the set of true nulls at level $\ell$ (including both exposed and non-exposed nulls), the total FWER across all levels is:
$\operatorname{FWER}\leq\sum_{\ell=1}^{L}\left(\left|\mathcal{H}_{0}^{(\ell)}\right|\cdot\alpha_{\ell}\cdot\prod_{j\in\text{anc}(\ell)}\theta_{j}\right)$ (B.13)
where $|\mathcal{H}_{0}^{(\ell)}|$ is the number of true nulls at level $\ell$ and $\text{anc}(\ell)$ is the set of ancestor nodes on the path from root to level $\ell$.
(c) Under the global null (the root is true, and therefore every node is true by the intersection structure), $\operatorname{FWER}\leq\alpha_{1}$. To see this: if the root is true, the stopping rule means no descendant is tested unless the root is first rejected. If the root is not rejected ($R_{1}=0$), no errors occur anywhere. If it is rejected ($R_{1}=1$), that rejection is itself a false positive. Either way, $\operatorname{FWER}=\mathbb{P}(R_{1}=1)\leq\alpha_{1}\leq\alpha$.
Proof.
For part (a), each exposed null at level $\ell$ has all non-null ancestors. The probability that all ancestors are rejected is $\prod_{j=1}^{\ell-1}\theta_{j}$ (the chain-rule product from Step 3 of the Proposition 1 proof). Conditional on reaching the node, the false rejection probability is at most $\alpha_{\ell}$ (by Valid Tests). The union bound says that the probability of at least one false rejection among the $e_{\ell}$ exposed nulls is at most the sum of their individual false-rejection probabilities — that is, $e_{\ell}\cdot\alpha_{\ell}\cdot\prod_{j=1}^{\ell-1}\theta_{j}$.
Part (b) follows by starting from the node-level bound in Proposition 1 and collecting all true nulls at the same level $\ell$. Each true null $i$ at level $\ell$ has test size $\alpha_{\ell}$ and an ancestor path product bounded by $\prod_{j\in\text{anc}(i)}\theta_{j}$. Summing first within each level and then across all levels yields the stated bound.
For part (c), if the root hypothesis is true — meaning there are no effects anywhere in the tree — then the intersection structure requires that every descendant is also true. No descendant is tested unless the root is rejected (by the Stopping Rule, Condition 3), and the root’s false rejection probability is at most $\alpha_{1}$ (by Valid Tests, Condition 4). Therefore $\operatorname{FWER}\leq\alpha_{1}$. ∎
Remark 5.
When tests at sibling nodes are conducted using independent randomizations (as they are in block-randomized experiments with independent blocks), the exact FWER contribution from $e_{\ell}$ exposed nulls at level $\ell$ is $\prod_{j=1}^{\ell-1}\theta_{j}\cdot[1-(1-\alpha_{\ell})^{e_{\ell}}]$. The union bound $e_{\ell}\cdot\alpha_{\ell}$ closely approximates $1-(1-\alpha_{\ell})^{e_{\ell}}$ when $\alpha_{\ell}$ is small. The union bound is simpler and sufficient for our purposes; we use it throughout.
Remark 6 (Where FWER Risk Concentrates).
Under the global null, FWER is exactly $\alpha_{1}\leq\alpha$ — the stopping rule at the root provides complete protection. The more interesting case arises when the root is false but some descendants are true nulls. If root power is high, the procedure almost always descends into the tree, exposing those true nulls to testing. This is precisely what Theorem 4(a) quantifies: the FWER contribution from level $\ell$ is the number of exposed nulls $e_{\ell}$, times the test size $\alpha_{\ell}$, times the path product to that level. When root power is high, the path product is close to 1, and the FWER contribution from each level is approximately $e_{\ell}\cdot\alpha_{\ell}$ — which can exceed $\alpha$ when many nulls are exposed. The next subsection formalizes the conditions under which this exposure remains manageable.
B.3 Natural Gating
In many practical settings—including the multi-site education policy trials that motivate this paper—the tree procedure controls strong FWER without any alpha adjustment. Splitting data across branches reduces sample size and therefore power at each level. When power decays fast enough relative to the branching factor, the procedure cannot reach enough true nulls to inflate the FWER. We formalize the relationship between power and error on the tree via the error load. In what follows, we will work with regular $k$-ary trees where each node has $k$ children. [^11] Note to readers of this draft: We have generalized this treatment in the code to the case of irregular $k$-ary trees where each node might have a different number of children and have verified FWER control using this concept using simulations on realistic data but have not yet added a formal extension to the theory here.
Definition 1 (Error Load).
For a regular $k$-ary tree with conditional rejection probabilities $\theta_{0},\theta_{1},\ldots,\theta_{L-1}$ along the root-to-leaf path, define the error load at level $\ell$ ($\ell=1,\ldots,L$) as
$G_{\ell}=k^{\ell-1}\prod_{j=0}^{\ell-1}\theta_{j},$ (B.14)
where $\theta_{0}=\theta_{\text{root}}$. The total error load is the sum of the level specific error loads, $\sum_{\ell=1}^{L}G_{\ell}$.
The error load $G_{\ell}$ measures the severity of exposure to false rejections at level $\ell$. It is the product of two quantities: $k^{\ell-1}$, the number of nodes at level $\ell$ in the full tree (the maximum number of true nulls that could be exposed there), and $\prod_{j=0}^{\ell-1}\theta_{j}$, the probability that the testing procedure reaches level $\ell$ by rejecting every ancestor along the way. When $G_{\ell}$ is small, either few nulls are exposed at that depth or the procedure is unlikely to reach it — either way, level $\ell$ contributes little to the FWER. When $G_{\ell}$ is large, many true-null nodes may be tested at that depth, and each such test has a false-rejection probability of at most $\alpha$ (by the Valid Tests condition). The total FWER contribution from level $\ell$ is therefore at most $G_{\ell}\cdot\alpha$.
Proposition 2 (Natural Gating Sufficiency).
If $\left(\sum_{\ell=1}^{L}G_{\ell}\right)\leq 1$ and each node is tested at level $\alpha$ (no adjustment), then the procedure controls FWER at level $\alpha$ for any configuration of true and false nulls:
$\mathrm{FWER}\leq\alpha.$
Proof.
By Proposition 1, we group exposed true nulls by level and apply the union bound within each level (the probability of at least one false rejection among $e_{\ell}$ exposed nulls is at most the sum of their individual false-rejection probabilities):
$\operatorname{FWER}\leq\sum_{\ell=1}^{L}\left(e_{\ell}\cdot\alpha\cdot\prod_{j=1}^{\ell-1}\theta_{j}\right).$ (B.15)
Each exposed null at level $\ell$ has all non-null ancestors, so $\prod_{j=1}^{\ell-1}\theta_{j}=\prod_{j=1}^{\ell-1}\pi_{j}$ (the product uses the power $\pi_{j}$ at each non-null ancestor). In a regular $k$-ary tree, the number of exposed nulls $e_{\ell}$ satisfies $e_{\ell}\leq k^{\ell-1}$ (the maximum when all siblings at level $\ell$ are null). Since $\theta_{j}\leq 1$ for all $j$, we can bound:
$\operatorname{FWER}\leq\alpha\cdot\left(\sum_{\ell=1}^{L}k^{\ell-1}\prod_{j=0}^{\ell-1}\theta_{j}\right)=\alpha\cdot\left(\sum_{\ell=1}^{L}G_{\ell}\right).$
When $\sum_{\ell=1}^{L}G_{\ell}\leq 1$, the quantity in parentheses is at most 1, so $\operatorname{FWER}\leq\alpha$. ∎
The key ratio governing whether the error load grows or shrinks with depth is:
$\frac{G_{\ell+1}}{G_{\ell}}=k\cdot\theta_{\ell}.$ (B.16)
When $\theta_{\ell}<1/k$, the error load decreases at level $\ell$; when $\theta_{\ell}>1/k$, it increases. Recall that $\theta_{\ell}$ is the conditional rejection probability (defined in (B.3)): it equals the power $\pi_{\ell}$ for non-null nodes and the test size $\alpha_{\ell}$ for true-null nodes. The ratio $k\cdot\theta_{\ell}$ compares the branching factor (how many new nodes appear at the next level) against the probability that each one is actually tested. When $\theta_{\ell}$ is below $1/k$, fewer nodes are tested at the next level than at the current level — the tree is “thinning out” as the procedure descends.
Corollary 1 (Critical Power Threshold).
If the conditional rejection probability at every level satisfies $\theta_{\ell}<1/k$, then $G_{\ell}$ decreases geometrically with depth. Writing $\theta_{\max}=\max_{\ell}\theta_{\ell}$ for the largest rejection probability at any level, the total error load satisfies $\sum_{\ell}G_{\ell}<G_{1}/(1-k\theta_{\max})$. Here $G_{1}=\theta_{\text{root}}$: when the root is false (as we expect in applications where we believe effects exist somewhere), $G_{1}$ equals the root power $\pi_{1}$. When $\theta_{\ell}<1/k$ at every level, this geometric sum converges and natural gating suffices to control the FWER.
The following corollary illustrates an extreme case: the root test has very low power, so the procedure almost never descends into the tree. This scenario is not desirable in practice — we want to detect effects when they exist — but it establishes a principle that will prove useful. The FWER depends on how far into the tree the procedure reaches, and low power at the root provides automatic FWER control by preventing the procedure from exposing any descendants to testing.
Corollary 2 (FWER Control via a Root-Level Bound).
Recall that the root node is indexed by $1$ and that the testing procedure above is used. Assume:
  1. 1.
    When $H_{1}$ is true, its conditional Type I error satisfies $\alpha_{1}\leq\alpha$. (This is easy to maintain for randomization-based tests and is the basis of the Valid Tests condition)
  2. 2.
    When $H_{1}$ is false, its conditional power satisfies $\pi_{1}\leq\alpha$. (That is, the test of the root has very low power.)
Then, for any configuration of true and false nulls in the tree,
$\mathrm{FWER}\leq\alpha.$
Proof.
This is the special case $G_{1}=\theta_{\text{root}}\leq\alpha<1$.
When $H_{1}$ is true, every node is true by the intersection structure, and $\operatorname{FWER}=\alpha_{1}\leq\alpha$.
When $H_{1}$ is false, every true null is a descendant of the root. A Type I error requires rejecting the root first:
$\operatorname{FWER}$ $=\mathbb{P}(R_{1}=1)\,\mathbb{P}(\text{at least one true descendant rejected}\mid R_{1}=1)$ (B.17)
$\leq\mathbb{P}(R_{1}=1)=\pi_{1}\leq\alpha.\qed$ (B.18)
Corollary 2 shows that low root power controls the FWER, but the mechanism — the procedure rarely reaches any descendants — is exactly the outcome we want to avoid. In practice, we design experiments to have enough power to detect effects when they exist. When root power is high, the procedure almost always descends into the tree, exposing true-null siblings to testing. We now examine how power decays with depth and then show what happens when root power is high.
Remark 7 (Connection to Data Splitting).
For a standardized effect size $\delta$ and total sample size $N$, the power at level $\ell$ when testing at significance level $\alpha$ is approximately
$\theta_{\ell}\approx\Phi\left(\delta\sqrt{\frac{N}{k^{\ell-1}}}-z_{1-\alpha/2}\right),$ (B.19)
where $\Phi$ is the standard normal CDF. Power decays because $n_{\ell}=N/k^{\ell-1}$, using equal splitting for simplicity. At the root, the full sample $N$ is available and power is typically high. At level $\ell$, the sample has been split among $k^{\ell-1}$ branches, so each branch uses roughly $N/k^{\ell-1}$ observations. Power drops because the test statistic’s signal-to-noise ratio scales with $\sqrt{n}$: splitting the sample among three branches cuts the effective sample by a factor of three and reduces the signal-to-noise ratio by $\sqrt{3}\approx 1.7$. In a tree with $k=3$ and moderate effect sizes, power typically drops below the critical threshold $1/k=1/3$ by level 3 or 4, so the error load decreases from that point onward.
In the multi-site education policy trials that motivate this paper, $k\approx 3$, $L\approx 3$, and power is moderate. Under these conditions, $\sum G_{\ell}<1$ and the unadjusted procedure controls FWER — explaining why the simulation in Section 3.3.1 of the main text shows $\mathrm{FWER}\leq\alpha$ without any adjustment.
Corollary 3 (FWER Inflation with High Root Power).
Suppose the root is false with root power $\pi_{1}=1$ (the root always rejects), and $e_{1}$ of the root’s $k$ children are true nulls ($1\leq e_{1}\leq k-1$; the intersection structure requires at least one child to be false). Since all nodes at level 2 are tested at the same significance level, which we write as $\alpha_{\text{lev.\,2}}$, then:
$\operatorname{FWER}=1-(1-\alpha_{\text{lev.\,2}})^{e_{1}}\approx e_{1}\cdot\alpha_{\text{lev.\,2}}\quad\text{for small }\alpha_{\text{lev.\,2}}.$ (B.20)
Without adjustment ($\alpha_{\text{lev.\,2}}=\alpha$), FWER exceeds $\alpha$ whenever $e_{1}\geq 2$, since $1-(1-\alpha)^{e_{1}}>\alpha$ for $e_{1}\geq 2$.
This corollary shows concretely what goes wrong when root power is high. When the root always rejects, the procedure tests every child of the root. If $e_{1}$ of those children are true nulls, each tested at level $\alpha$, then the probability of at least one false rejection is $1-(1-\alpha)^{e_{1}}$, which grows roughly in proportion to $e_{1}$.
Example. With $k=5$ branches and $e_{1}=4$ true-null children (one child is non-null), each tested at $\alpha=0.05$: $\text{FWER}=1-0.95^{4}=0.186$, nearly four times the nominal rate. With $k=10$ and $e_{1}=9$: $\text{FWER}=1-0.95^{9}=0.370$. The adaptive adjustment of the next subsection corrects this by testing each child at $\alpha/k$ rather than $\alpha$, so that the total stays bounded.
B.4 Adaptive Alpha Adjustment When Gating Is Insufficient
When $\sum_{\ell}G_{\ell}>1$ — because the tree is wide, deep, and root power is high — the unadjusted procedure inflates FWER. The adaptive $\alpha$-adjustment divides the nominal $\alpha$ by the error load at each level, compensating for the number of true-null nodes the procedure can reach at that depth. This ensures that the total FWER stays bounded.
Remark 8 (Adaptive Alpha Adjustment with Power Decay on Regular $k$-ary Trees).
The adjusted significance levels at level $\ell$ are:
$\alpha_{\ell}^{\text{adj}}=\min\left\{\alpha,\frac{\alpha}{k^{\ell-1}\cdot\prod_{j=1}^{\ell-1}\hat{\theta}_{j}}\right\}$ (B.21)
where $\hat{\theta}_{j}$ is the estimated conditional rejection probability at level $j$, computed from the power decay model (B.19) using an estimated effect size $\hat{\delta}$.
Algorithm 1 Adaptive Alpha Adjustment Accounting for Power Decay
0: Tree structure $(k,L)$, sample size $N$, nominal $\alpha$, effect size estimate $\hat{\delta}$
0: Adjusted significance levels $\{\alpha_{\ell}^{\text{adj}}\}_{\ell=1}^{L}$
1: Calculate estimated power at each level:
2:for $\ell=1$ to $L$ do
3:  $n_{\ell}\leftarrow N/k^{\ell-1}$
4:  $\hat{\theta}_{\ell}\leftarrow\Phi\left(\hat{\delta}\sqrt{n_{\ell}}-z_{1-\alpha/2}\right)$
5:end for
6: Calculate error load: $G_{\ell}\leftarrow k^{\ell-1}\prod_{j=1}^{\ell}\hat{\theta}_{j}$
7:if $\sum_{\ell=1}^{L}G_{\ell}\leq 1$ then
8:  Natural gating sufficient: $\alpha_{\ell}^{\text{adj}}\leftarrow\alpha$ for all $\ell$
9:else
10:  for $\ell=1$ to $L$ do
11:   if $\ell=1$ then
12:    $\alpha_{1}^{\text{adj}}\leftarrow\alpha$ {Root level}
13:   else
14:    $\text{PathPower}\leftarrow\prod_{j=1}^{\ell-1}\hat{\theta}_{j}$
15:    $\alpha_{\ell}^{\text{adj}}\leftarrow\frac{\alpha}{k^{\ell-1}\times\text{PathPower}}$
16:   end if
17:  end for
18:end if
19:return $\{\alpha_{\ell}^{\text{adj}}\}_{\ell=1}^{L}$
Remark 9 ($p$-value monotonicity as algorithmic coherence).
One could additionally require that $p_{\text{child}}\geq p_{\text{parent}}$ at every node, enforcing this by replacing each child $p$-value with $\max(p_{\text{child}},p_{\text{parent}})$. This would guarantee that rejection decisions are coherent with the tree structure: if a parent is not rejected, no child would be rejected either, so the Stopping Rule (Condition 3) would be satisfied automatically by the $p$-values rather than imposed by the algorithm. Monotonicity also ensures that power decays with depth, aligning the algorithm’s behavior with the power-decay model in (B.19). However, none of the results in this section require $p$-value monotonicity: the FWER bounds in Proposition 1, Theorem 4, and Theorem 9 below hold for any valid tests satisfying the Stopping Rule. The manytestsr R package does not enforce $p$-value monotonicity, and the simulations in this paper confirm that control holds without monotonicity: the weak FWER simulations yield identical results with and without it, and the strong FWER simulations (Section 3.3.1) draw parent and child $p$-values independently.
{theorem}
[FWER Control Under Adaptive Adjustment] Algorithm 1 controls FWER at level $\alpha$ for any configuration of null hypotheses, provided the power estimates $\hat{\theta}_{\ell}$ are not underestimated (i.e., $\hat{\theta}_{j}\geq\theta_{j}$ for all non-null ancestors $j$).
The intuition: at each level of the tree, the number of true-null nodes actually tested depends on how many non-null ancestors were rejected. The product $\prod\hat{\theta}_{j}$ estimates the probability that the procedure descends to a given depth. If this descent probability is high, many true nulls may be exposed to testing, so we compensate with a more stringent $\alpha_{\ell}^{\text{adj}}$. The adaptive formula divides the nominal $\alpha$ by the expected number of exposed true nulls at each level.
Proof.
We show $\operatorname{FWER}\leq\alpha$ for any configuration of true and false nulls, assuming $\hat{\theta}_{j}\geq\theta_{j}$ for all non-null ancestors (power is not underestimated).
Setup. Call a true null node exposed if its parent is non-null (or if it is the root and is null). Let $m_{\ell}$ be the number of non-null nodes at level $\ell$, and $e_{\ell}$ the number of exposed nulls at level $\ell$. In a regular $k$-ary tree, each non-null parent at level $\ell-1$ has exactly $k$ children. Those $k\cdot m_{\ell-1}$ children fall into two categories: some are themselves non-null ($m_{\ell}$ of them), and the rest are exposed true nulls ($e_{\ell}$ of them — “exposed” because their non-null parent will likely be rejected, opening them to testing). This gives the identity:
$m_{\ell}+e_{\ell}=k\cdot m_{\ell-1},\qquad\ell=2,\ldots,L.$
Contribution of an exposed null at level $\ell$. An exposed null at level $\ell$ has all non-null ancestors. With the adaptive adjustment $\alpha_{\ell}^{\text{adj}}=\alpha/(k^{\ell-1}\prod_{j=1}^{\ell-1}\hat{\theta}_{j})$, this node’s contribution to the FWER upper bound is:
$\underbrace{\frac{\alpha}{k^{\ell-1}\prod_{j=1}^{\ell-1}\hat{\theta}_{j}}}_{\text{adjusted test size }\alpha_{\ell}^{\text{adj}}}\cdot\underbrace{\prod_{j=1}^{\ell-1}\pi_{j}}_{\text{probability of reaching level }\ell}\leq\frac{\alpha}{k^{\ell-1}},$
where the inequality uses $\hat{\theta}_{j}\geq\pi_{j}$ for non-null ancestors (the conservative-direction assumption).
Contribution of non-exposed nulls. A non-exposed null at level $\ell^{\prime}$ descends from a true-null ancestor at some level $\ell<\ell^{\prime}$. To reach this non-exposed null, the procedure must falsely reject that true-null ancestor — an event with probability at most $\alpha_{\ell}^{\text{adj}}\leq\alpha$. This means the path product for a non-exposed null contains at least one factor of $\alpha$ (from the null ancestor) in addition to the $\alpha/k^{\ell^{\prime}-1}$ from the node’s own test size. The total contribution is therefore at most $\alpha^{2}/k^{\ell^{\prime}-1}$: a product of two small numbers. For $\alpha=0.05$, this is at most $0.0025/k^{\ell^{\prime}-1}$, which is negligible compared to the exposed-null contributions of $\alpha/k^{\ell-1}$. We can safely ignore these terms in what follows.
The telescoping sum. The following manipulation is called a “telescoping sum” because most terms cancel in pairs, like the sections of a collapsing telescope. We substitute the identity $e_{\ell}=k\cdot m_{\ell-1}-m_{\ell}$ (from the Setup above), split the sum into two parts, and find that all intermediate terms cancel, leaving only the first and last.
Summing over exposed nulls:
$\operatorname{FWER}\leq\alpha\sum_{\ell=2}^{L}\frac{e_{\ell}}{k^{\ell-1}}+\text{(non-exposed terms)}.$
Substituting $e_{\ell}=k\cdot m_{\ell-1}-m_{\ell}$:
$\sum_{\ell=2}^{L}\frac{e_{\ell}}{k^{\ell-1}}$ $=\sum_{\ell=2}^{L}\frac{k\cdot m_{\ell-1}-m_{\ell}}{k^{\ell-1}}=\sum_{\ell=2}^{L}\frac{m_{\ell-1}}{k^{\ell-2}}-\sum_{\ell=2}^{L}\frac{m_{\ell}}{k^{\ell-1}}.$
Re-index the first sum with $s=\ell-1$ (so $s$ runs from $1$ to L1L-1):
$=\sum_{s=1}^{L-1}\frac{m_{s}}{k^{s-1}}-\sum_{s=2}^{L}\frac{m_{s}}{k^{s-1}}=\frac{m_{1}}{k^{0}}-\frac{m_{L}}{k^{L-1}}=m_{1}-\frac{m_{L}}{k^{L-1}}.$
Since $m_{1}\leq 1$ (the root is at most one node) and $m_{L}\geq 0$:
$m_{1}-\frac{m_{L}}{k^{L-1}}\leq 1.$
The telescoping sum shows that $\sum_{\ell=2}^{L}e_{\ell}/k^{\ell-1}\leq m_{1}-m_{L}/k^{L-1}$. Since $m_{1}$ is the number of non-null nodes at the root level, and there is only one root, $m_{1}\leq 1$. And $m_{L}/k^{L-1}\geq 0$ because it is a count divided by a positive number. Subtracting a non-negative quantity from something at most 1 gives something at most 1. Substituting back into the FWER bound: the exposed-null contribution is at most $\alpha$ times this sum, which is at most $\alpha\cdot 1=\alpha$. The non-exposed terms (shown above to be at most $\alpha^{2}/k^{\ell^{\prime}-1}$ each) are negligible by comparison and do not change the bound. Therefore $\operatorname{FWER}\leq\alpha$.
If the root is itself a true null ($m_{1}=0$), then the global null holds: every node in the tree is true. In this case, the stopping rule provides complete protection — the FWER equals $\mathbb{P}(R_{1}=1)\leq\alpha_{1}\leq\alpha$, because rejecting the root is the only way any error can occur. The telescoping sum gives $\sum e_{\ell}/k^{\ell-1}=0-m_{L}/k^{L-1}\leq 0$, which is non-positive, so the bound is automatically satisfied. ∎
Remark 10 (The $\tau$-relaxation).
In practice, one may wish to relax the adaptive adjustment at deep levels where power is already very low. Algorithm 1 includes a natural gating check ($\sum G_{\ell}\leq 1$) as the first step. When this check passes, no adjustment is needed at any level. When it fails, one could in principle relax the adjustment at levels where the remaining error load $\sum_{\ell^{\prime}\geq\ell}G_{\ell^{\prime}}$ is small. This is where natural gating “takes over” from the adaptive adjustment.
The formal guarantee of Theorem 9, however, requires the full adjustment at every level. Relaxing the adjustment at levels where power is low but the error load has not yet been spent can inflate FWER. For example, with $k=10$, $L=2$, and a false root with high power, 10 null children each tested at nominal $\alpha=0.05$ give $\mathrm{FWER}\approx 1-0.95^{10}\approx 0.40$. Users who wish to relax the adjustment beyond the natural gating check should verify FWER control empirically via simulation.
Appendix C Test statistics with power against diverse alternatives.
This approach requires, in [undeft]’s words, a “first true null hypothesis”, or an ordering of hypotheses such that non-rejection of the first hypothesis ought to imply a non-rejection of the subsequent hypotheses. This requirement raises questions about the test statistic used. A difference-of-means test statistic has power to detect shifts in the mean. But such statistics lose power in the presence of outliers [undefad], so a test on a smaller subset without outliers could have more power than a test on the full sample. Moreover, mean-difference statistics record the sign of effects: if half the blocks have positive effects and half negative, the overall mean difference may be near zero despite genuine effects in every block. In either case, the root-level test could fail to reject even when block-level effects are real, preventing the procedure from descending into the tree.
C.1 A Combined Energy Statistic
We introduce a test statistic that captures not only mean differences but also distributional differences between treated and control units, in ways that are not fooled by strong countervailing effects. The strategy is to transform each unit’s outcome into several representations — ranks, pairwise distances [undefy], and nonlinear transforms — and then combine them into a single test. Specifically, we compute the following six scores for each unit:
  1. 1.
    The raw outcome
  2. 2.
    The rank of the outcome
  3. 3.
    The mean Euclidean distance between a unit’s outcome and the outcomes of all other units in the block, following the energy-distance framework of [undefz, undefy, undefs].
  4. 4.
    The mean Euclidean distance between the unit’s outcome rank and the ranks of all other units.
  5. 5.
    The maximum Euclidean distance between the unit and all other units.
  6. 6.
    The hyperbolic tangent (tanh) transform of the outcome, which serves as an alternative to the log transform when the variable has many zeros [undefag].
Each unit thus receives six scores, each capturing a different aspect of the outcome distribution. We combine these scores into a single test statistic and compare the observed value to its randomization distribution, using either large-sample chi-square approximations [undefg, undefaf, undefx] or exact permutation tests when the sample size is small.
Why should this somewhat arbitrary combined test statistic have large-sample approximations? Under the null hypothesis of no effects and given the known randomization (within blocks in this case), a linear combination of mean differences follows a chi-square distribution. The required variances and covariances are functions of the randomization distribution, which the null and the design determine [undefg, undefx]. The independence_test function in the coin R package [undefl] implements this combination.
We do not claim that this omnibus test is optimal in all circumstances. But a tree-structured testing procedure needs a root-level test that is sensitive to diverse departures from the null, and the combined energy statistic meets this requirement without excessive computational cost.
C.2 The Combined Stephenson Rank Test
[undefm] present a test that targets the distribution of treatment effects across units in a block-randomized experiment. Specifically, it tests hypotheses about the additive effect $c$ at each quantile of the experimental pool, answering questions such as “what proportion of units benefited from the intervention?” If that proportion exceeds zero, at least one unit benefited — which counts as a detected effect in our framework. We apply this test via the CMRSS R package of [undefm].
Appendix D MDRC Application: Full Results
Table D.2 reports the distribution of the outcome variable (credits earned in the first main session after the intervention) across all 25 studies.
Study Blocks min med mean max
CUNY Start 21 0 0.00 2.28 27
ASAP Ohio 9 0 10.00 9.19 32
OD PBS + Advising 11 0 6.00 5.97 27
ASAP CUNY 5 0 11.75 10.36 31
EASE 26 0 0.00 1.42 21
LC Career 28 0 12.00 11.24 26
OD LC 4 0 12.00 10.96 26
PBS OH 11 0 8.00 8.09 29
OD Success 2 0 3.00 3.92 17
ModMath 4 0 6.00 6.01 26
DPP 44 0 0.00 4.33 16
PBS + Supports 15 0 9.00 8.92 43
LC English 4 0 6.00 6.18 30
PBS + Math 6 0 8.00 8.07 34
iPASS MCCC 42 0 6.00 5.52 28
DCMP 31 0 6.00 6.55 27
OD Advising + Incentive 8 0 6.00 5.80 30
iPASS UNCC 35 0 14.00 13.14 30
PBS Variations 6 0 0.00 4.02 29
iPASS Fresno State 8 0 12.00 11.84 27
AtD Success Course 8 0 7.00 6.80 25
LC Reading 7 0 7.00 7.28 29
PBS + Advising 2 0 14.00 12.92 21
OD Success (Enhanced) 2 0 1.00 3.08 17
LC English + Success 4 0 6.50 6.13 23
Table D.2: Distribution of the outcome variable (credits earned in the first main session after the intervention) across all 25 MDRC studies. The outcome is skewed toward 0: in most studies the median student earns few credits. See Table 5 for the testing results.
Table D.3 reports the structured testing results for all 25 MDRC studies, including the 13 studies where neither overall test rejected at $\alpha=0.05$. Table 5 in the main text presents the 12 studies with at least one significant overall test.
[…]
Table D.3: Structured (top-down) vs. bottom-up testing in all 25 MDRC studies. See Table 5 in the main text for the subset with at least one significant overall test. ’3 Rules’ applies the stopping rule and valid tests only. ’+ Adj $\alpha$ Pr.’ adds the adaptive $\alpha$-adjustment with branch pruning. ’Ratio’ is top-down node detections divided by bottom-up Hommel detections.